# AuthGlass: Enhancing Voice Authentication on Smart Glasses via Air-Bone Acoustic Features

### v1

@@ 6.4. Limited Sampling Rate
6.4.1. Method
Another aspect that limit the real-world usage was the sampling rate, as our prototype samples audio at a frequency of 96 kHz, requiring relatively high energy and memory demands. In contrast, most commercial smart glasses adopt microphones with a 16 kHz sampling rate. While such a rate has little effect on Mel-spectrogram computation, it can impact sound field modeling and time-delay estimation accuracy. To verify the effect of limited sampling rates on voice authentication, we down-sampled the acoustic data to 16 kHz, and re-tested the authentication EER, with other configurations same as those in section 4.5.
6.4.2. Result and Discussion
As shown in Table 7, our model continues to achieve strong performance at a lower sampling rate of 16kHz. The results indicate that our model performs well across different sampling rates, demonstrating robustness to variations in input resolution.
AC AC+DMs AC+BC AC+BC+DMs
mean EER 0.0917 0.0460 0.0648 0.0302
Table 7. Averaged EER of Different Feature Combination for 16kHz Sampling Rate.
@@ 7. Discussion
7.1. Conclusion of Findings
Using our prototype, we observed that both vibration and sound-field features are closely related to human speech production. To extract these features, we applied the well-established log Mel-spectrogram method for vibration signals, and we further proposed two novel representations—time delay matrix (TDM) and energy decrease matrix (EDM)—to capture temporal variations in sound-field features. With a relatively simple authentication architecture, we demonstrated that vibration and sound-field features each improve authentication performance, and their combination yields even better results. Moreover, extensive experiments showed that the proposed AuthGlass system can effectively resist attacks while maintaining robust performance under practical conditions.
@@ 7.3.3. Robustness against Movement Noises
7.3.4. Deployment to commercial devices
As one of the key focuses of this work is the practical applicability of AuthGlass, several challenges remain for deployment on commercial devices. Although the TDM and EDM features proposed in this paper have been validated as effective, their computation still incurs a non-negligible processing time, which limits the feasibility of real-time authentication. In addition, the system relies on precise synchronization of the microphone arrays to accurately capture sound-field data, posing additional design and implementation challenges. Furthermore, there is currently no mature solution for embedding bone-conduction microphones into the nosepad of smart glasses, which may restrict both signal quality and overall system performance. Despite these limitations, advances in hardware and embedded systems present opportunities to overcome these challenges. Improved sensor integration, more efficient signal processing algorithms, and faster computation could not only resolve current hardware constraints but also enhance the real-time capability and robustness of AuthGlass, paving the way for its adoption in practical, commercial applications.

### v2

@@ 6.4. Limited Sampling Rate
Another aspect that limit the real-world usage was the sampling rate, as our prototype samples audio at a frequency of 96 kHz, requiring relatively high energy and memory demands. In contrast, most commercial smart glasses adopt microphones with a 16 kHz sampling rate. While such a rate has little effect on Mel-spectrogram computation, it can impact sound field modeling and time-delay estimation accuracy. To verify the effect of limited sampling rates on voice authentication, we down-sampled the acoustic data to 16 kHz, and re-tested the authentication EER, with other configurations same as those in section 4.3. The results indicate that our model performs well across different sampling rates, demonstrating robustness to variations in input resolution.